OpenStudy (anonymous):

Using L'Hopital Ruleto evaluate 5x(cos 9x-1)/sin 8x-8x

OpenStudy (anonymous):

Take the derivative on both numerator and denominator...

OpenStudy (anonymous):

Yes i get down to the third derivative and I jsut can not seem to get the answer correct. here 5limx app. 0 [9x sin-5cos9x] / 8 limx app. 0 [-8cos8x]

OpenStudy (anonymous):

That is the third derivative

OpenStudy (anonymous):

Show me how you derivated the numerator? Show me the derivative of : x(cos9x - 1) Show me..

OpenStudy (anonymous):

5limx app0 [-9xsin 9x+ cos 9x-1]

OpenStudy (anonymous):

Yes now you are doing right..

OpenStudy (anonymous):

Now do the derivative of : sin8x - 8x..

OpenStudy (anonymous):

now that equals to 0/0making that derivative indeterminate right so i kept going but i get stuck on the last part

OpenStudy (anonymous):

lim x app 0 [8 cos 8x]

OpenStudy (anonymous):

You cannot take out 8 as common of it.. sin8x here 8x is an angle -8x it is just a value..

OpenStudy (anonymous):

Do the derivative separately for them: Tell me what is the derivative of sin8x?

OpenStudy (anonymous):

so [ 8 cos 8x-8] is incorrect

OpenStudy (anonymous):

Now you are right..

OpenStudy (anonymous):

okay i made an errror in typing it in

OpenStudy (anonymous):

Now apply the limits.. cos0 = 1 sin0 = 0 Use this to get the answer..

OpenStudy (anonymous):

0/0

OpenStudy (anonymous):

No you are incorrect..

OpenStudy (anonymous):

Numerator is 0 no doubt but denominator has a certain value.. Check it once more..

OpenStudy (anonymous):

0/8

OpenStudy (anonymous):

No no sorry you are right.. Now take the derivative once more...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the next derivative is numerator

OpenStudy (anonymous):

Yeah tell me the value after derivative..

OpenStudy (anonymous):

is 5 lim x app. 0 [-18x cos 9x -18 sin9x] / 4 lim x app 0 [-8 sin 8x] = 0/0

OpenStudy (anonymous):

It is big so take your time and do not do it hastly go slowly I can wait..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9*9 = 81 and not 18..

OpenStudy (anonymous):

ok sorry yeah error again 5 lim x app. 0[-81x cos 9x -81 sin9x]

OpenStudy (anonymous):

Just and idea. After taking the derivative the first time, try to split into 2 fractions. Do so after taking the derivative the second time.

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

[-9xsin 9x+ cos 9x-1] = x(-81cos9x) + (-9sin9x) - 9sin9x = -81xcos9x - 18sin9x I think I got this much the numerator.. Isn't it??

OpenStudy (anonymous):

oh is it all wrong

OpenStudy (anonymous):

I am wrong or you are wrong??

OpenStudy (anonymous):

oh what did you get

OpenStudy (anonymous):

-81xcos9x - 18sin9x Getting it or not?

OpenStudy (anonymous):

i am confuesed this is the part i tend to mess up maybe i am wrong i do not know

OpenStudy (anonymous):

See we are going at very fast rate.. Let us do it slowly.. Find the derivative of (-9xsin9x) first and tell me what you got?

OpenStudy (anonymous):

ok yeah its just that i have it worked out but let me work it out

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

-81x cos 9x

OpenStudy (anonymous):

You have to use product rule here @blopez ... So you will get one more term with it..

OpenStudy (anonymous):

\[\frac{d}{dx}uv = u.\frac{dv}{dx} + v.\frac{du}{dx}\]

OpenStudy (anonymous):

I got the answer. 1215/512

OpenStudy (anonymous):

yes okay -81x cos 9x-81

OpenStudy (anonymous):

i need help working the problem out not the answer but thanks

OpenStudy (anonymous):

@Mr._To thanks for the answer.. I will carry blopez to that answer...

OpenStudy (anonymous):

If you can wait for 5 minutes, I'll post my work.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How you got -81?? Just use the formula and check for answer once more how you did it..

OpenStudy (anonymous):

so next is sin 9x

OpenStudy (anonymous):

because -9*9

OpenStudy (anonymous):

or does the sign change because from sin to cos?

OpenStudy (anonymous):

i do not understand why it is not [-81x cos 9x -81 sin9x] if i am using the product rule or am i missing something

OpenStudy (anonymous):

Tell me what is the value of u? Compare (-9x)(sin9x) with uv and tell what is u from here?

OpenStudy (anonymous):

See, I will show you in more clearer form: Wait a minute..

OpenStudy (anonymous):

-9x is u

OpenStudy (anonymous):

And so v is sin9x and now use the formula slowly and solve for u and v..

OpenStudy (anonymous):

If you need steps in between, please let me know.

OpenStudy (anonymous):

-81 x cos 9x

OpenStudy (anonymous):

\[\frac{d}{dx}(-9x)(\sin9x) = (-9x)\frac{d(\sin9x)}{dx} + \sin(9x).\frac{d(-9x)}{dx}\] Can you now solve this??

OpenStudy (anonymous):

You are finding the first term only but you have to solve for other term also.. See I have written above in more understandable form Please solve the above one..

OpenStudy (anonymous):

[-81x cos 9x - 18 sinx 9x]

OpenStudy (anonymous):

Now you got it.. Getting how it came??

OpenStudy (anonymous):

yes sorry it took so long

OpenStudy (anonymous):

Now we are left with the other terms in the numerator that are : Cos9x - 1 Can you find the derivative of these two? tell me what is that.. No need to say sorry I am just helping you and I will remain helping you until you did get it..

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Solving it further or not??

OpenStudy (anonymous):

yes iam

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

sin 9x ??

OpenStudy (anonymous):

What is derivative of cosx? I think it is -sinx not sinx Be careful @blopez

OpenStudy (anonymous):

yes it is so it is -sin 9x

OpenStudy (anonymous):

\[\frac{d}{dx}\cos(ax) = -a.\sin(ax)\] Use the above formula...

OpenStudy (anonymous):

It is not -sin9x there is one more mistake in this..

OpenStudy (anonymous):

ok -9 sin 9x

OpenStudy (anonymous):

Yes.. We found them separately now join all the derivatives to get numerator.. What it came??

OpenStudy (anonymous):

here it goes 5 lim x app. 0 [-81x cos 9x-18 sin 9x]

OpenStudy (anonymous):

Yes now take the derivative of denominator...

OpenStudy (anonymous):

ok 4 lim x app. 0 [-8 cos 8x]

OpenStudy (anonymous):

Firstly tell me what was the denominator before derivative??

OpenStudy (anonymous):

ok [cos 8x-1]

OpenStudy (anonymous):

Have you taken 8 out of limits?

OpenStudy (anonymous):

8 lim x app. 0 [cos 8x-1]

OpenStudy (anonymous):

Limit x app 0 (8cos8x - 8) This is the denominator or you can take it as: 8Limit x app 0 (cos8x - 1).. Getting or not?

OpenStudy (anonymous):

Now tell me the derivative of cos8x??

OpenStudy (anonymous):

-8sin 8x or just - 8 sin

OpenStudy (anonymous):

-8 sin 8x

OpenStudy (anonymous):

Yes now you got it..

OpenStudy (anonymous):

So now we are left with last derivative ie the third derivative.. So, be calm and take the third derivative of the numerator.. Let us do it...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Firstly slowly take the derivative of (-81x)(cos9x) \[\frac{d}{dx}(-81x)(\cos9x) = (-81x).\frac{d(\cos9x)}{dx} + (\cos9x).\frac{d(-81x)}{dx}\] Solve for this...

OpenStudy (anonymous):

ok this is what i had worked out before coming here for the numerator 5 lim x app. 0 d/dx [-x cos 9x -sin 9x] 5 limx app. 0[ 9x sin 9x-5 cos 9x]

OpenStudy (anonymous):

How you did this??

OpenStudy (anonymous):

is it wrong.

OpenStudy (anonymous):

Yes totally... Just go as I am saying otherwise you will confuse yourself..

OpenStudy (anonymous):

ok jaja

OpenStudy (anonymous):

Jaja means??